find points of intersection of

f(x)=x^4+8x^3+16x^2

g(x)=5-4x^2-16x


1
Expert's answer
2022-01-27T01:00:16-0500

"f(x)=x^4+8x^3+16x^2"

"g(x)=5-4x^2-16x"

To find the point of intersection, let's equate both equations. This gives

"x^4+8x^3+16x^2=5-4x^2-16x"

"x^4+8x^3+16x^2-5+4x^2+16x=0"

"x^4+8x^3+16x^2+4x^2+16x-5=0"

"x^4+8x^3+20x^2+16x-5=0"

"x^4+" "8x^3+20x^2+16x=5"

"x(x+2)^2(x+4)=5"

"\left[x(x+4)-1\right]\left[x(x+4)+5\right]=0"

"[x^2+4x-1][x^2+4x+5]=0"


Hence, we set "[x^2+4x-1]=0\hspace{0.2cm}and\hspace{0.2cm}[x^2+4x+5]=0"

"[x^2+4x-1]=0"

solving quadratically, we have

"\displaystyle \implies x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\\\displaystyle \implies x=\frac{-4\pm\sqrt{4^2-4(1)(-1)}}{2(1)}"


"\displaystyle \implies\frac{-4\pm\sqrt{16+4}}{2}"


"\displaystyle \implies\frac{-4\pm\sqrt{20}}{2}"


"\displaystyle \implies\frac{-4}{2}\pm\frac{\sqrt{20}}{2}"


"\displaystyle \implies-2\pm\frac{2\sqrt{5}}{2}"


"\displaystyle \implies-2\pm{\sqrt{5}}"


"\displaystyle \implies-2+{\sqrt{5}}\hspace{0.2cm}or\hspace{0.2cm}\displaystyle \implies-2-{\sqrt{5}}"


Also, solving the second Quadratic equation similarly, we have

"[x^2+4x+5]=0"

"\displaystyle \implies x=\frac{-4\pm\sqrt{4^2-4(1)(5)}}{2(1)}"


"\displaystyle \implies\frac{-4\pm\sqrt{16-20}}{2}"


"\displaystyle \implies\frac{-4\pm\sqrt{-4}}{2}"


"\displaystyle \implies\frac{-4}{2}\pm\frac{\sqrt{-4}}{2}"


"\displaystyle \implies-2\pm\frac{2i}{2}"


"\displaystyle \implies-2\pm i"

"\implies -2+i\hspace{0.2cm}or\hspace{0.2cm}-2-i"

Finding the roots of the polynomial,

"x=(-\sqrt5-2,-2+\sqrt5,-2-i,-2+i)"

let "y=f(x)=g(x)"

substituting "x" into either "f(x) or g(x)" we get the same value for "y."

Let's substitute "x" into "g(x)"

"y=5-4x^2-16x"

"\implies y=(1,25)"

The points of intersection are

"(-\sqrt5-2,1),(-2+\sqrt5,1),(-2-i,25),(-2+i,25)."